# Covariate-Adaptive Randomization in Clinical Trials without Inflated Variances

### v2

@@ 4.1 The properties of the CAR procedures
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+O\Big(n^{\frac{\gamma}{2}}+n^{\frac{1}{4}+\frac{1}{4(r_{0}-1)}}+n^{1-\gamma}+n^{\frac{3}{2}+\frac{1}{2(r_{0}-1)}-2\gamma}\Big)$
$=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+o(n^{1/2})\text{ in }L_{1}$
by the facts that $\frac{\gamma}{2}<\frac{1}{2}$, $\frac{1}{4}+\frac{1}{4(r_{0}-1)}<\frac{1}{2}$, $1-\gamma<\frac{1}{2}$ and $\frac{3}{2}+\frac{1}{2(r_{0}-1)}-2\gamma<\frac{1}{2}$.
Notice that $\{(\Delta M_{i,\tilde{Z}},W_{i}),\mathscr{F}_{i},i=1,2,\ldots\}$ is sequence of martingale differences. We have $|\Delta M_{i,\tilde{Z}}|\leq|\widetilde{Z}_{i}|+\textsf{E}|\widetilde{Z}_{i}|$,
@@
i.e., the Linderberg condition is satisfied. Further, and $\textsf{E}[W_{i}^{2}|\mathscr{F}_{i-1}]=\textsf{E}W^{2}$,
@@
$\left(\frac{M_{n,\widetilde{Z}}}{\sqrt{n}},\frac{\sum_{i=1}^{n}W_{i}}{\sqrt{n}}\right)\overset{d}{\to}N(0,\sigma^{2})\times N(0,\textsf{Var}(W)),$
by the central limit theorem for martingales(Hall and Heyde, , 1980, Corollary 3.1). Hence, (2.8) holds. $\Box$
Proof of Theorem 2.2. We will use the induction method to obtain the moment estimates of $\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]$ by inducting both in $n$ and $r$. We assume
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{s}]\leq C_{s}n^{\beta s/2}\text{ for all integers }n,s\geq 0,$ (4.40)
and
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r-2}]\leq C_{r}n^{\alpha(r-2)/2}\text{ for all }n,$ (4.41)
where $0<\beta/2<\alpha\leq\beta\leq 2$.
Let $q$ be an integer large enough such that
$\frac{(\beta-\alpha)r-\beta}{2q}+\frac{\beta}{2}-\alpha<0.$
Then by (4.1),
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq$ $C_{r}\left(m^{\frac{\alpha(r-2)}{2}+\gamma}+m^{\frac{\alpha(r-2)}{2p}+\frac{\beta(r-1+q)}{2q}}\right)$
$\leq$ $C_{r}\left(n^{\frac{\alpha r}{2}+\gamma-\alpha}+n^{\frac{\alpha r}{2}+\frac{(\beta-\alpha)r-\beta}{2q}+\frac{\beta}{2}-\alpha}\right)$
$\leq$ $C_{r}n^{\alpha r/2}\;\;\text{ for all }n,$ (4.42)
whenever $\alpha\geq\gamma$.
Obviously, (4.40) holds for $\beta=1$ by (4.34), and (4.41) is true for $r=2$. So, by (4.1) and the induction, (4.41) holds for all integer $r\geq 2$ whenever $\alpha\geq\gamma$ and $\alpha>1/2$, which implies that (4.40) holds for any $\beta$ with $\beta\geq\gamma$ and $\beta>1/2$. Repeating the above procedure $i$ times concludes that (4.40) holds for any $\beta$ with $\beta\geq\gamma$ and $\beta>1/2^{i-1}$. Stooping the procedure when $1/2^{i-1}<\gamma$, we conclude that
$\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq C_{r}n^{\gamma r/2}\;\text{ for all integers }n\text{ and }r\geq 2.$
(2.9) is proven.
When $\textsf{E}|Z|<\infty$, the proofs of (2.6) and (2.7) are the same as those in Theorem 2.1 since $\textsf{E}\|\bm{\Lambda}_{n}\|=o(n^{\gamma})$.
Suppose $\textsf{E}[Z^{2}]<\infty$. Then ((ii)) follows from (2.9) and (4.38), and, (2.8) follows from ((ii)) and (4.38). Notice
$\textsf{E}M_{n,\widetilde{Z}}^{2}=\sum_{i=1}^{n}\textsf{E}(\Delta M_{i,\tilde{Z}})^{2}\leq n\textsf{E}\widetilde{Z}^{2}\leq n\textsf{E}Z^{2}.$
By (4.38) and (2.9), (2.12) holds. $\Box$
Proof of Remark 2.2. Notice
$\left|\ell(x)-\rho-\ell^{\prime}(0)x-\frac{1}{2}\ell^{\prime\prime}(0)x^{2}\right|\leq c|x|^{3}.$
Similar to (4.1), we have
@@
$\leq$ $C\textsf{E}\left[\frac{\|\bm{\Lambda}_{n}\|^{3}\|\bm{\Phi}\|^{3}|\widetilde{Z}|]}{n^{2\gamma}}\big|\bm{\Lambda}_{n}\right]\leq C_{0}(\textsf{E}Z^{2})^{1/2}\frac{\|\bm{\Lambda}_{n}\|^{3}}{n^{3\gamma}}.$
@@
$=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+O(n^{\gamma/2}+n^{1-3\gamma/2})\text{ in }L_{r},$
where $C_{0}$ does not depend on $Z$.
Proof of Theorem 2.3. Notice $\Phi(u_{\rho/2})<\rho<\Phi(u_{(\rho+1)/2})$. Let $K>2$. By the continuity, there exists $\delta>0$ such that
@@
When $|x|\geq\delta$, we have $|\ell(x)-\rho|\leq 2\leq 2\delta^{-K}|x|^{K}$, $|x|\leq\delta^{-K+1}|x|^{K}$. So, there exists a constant $C_{K}$ such that
@@
$\;\;\leq c_{K}\textsf{E}\left[\frac{\|\bm{\Lambda}_{n}\|^{K}\|\bm{\Phi}\|^{K}|Z|]}{n^{2\gamma}}\big|\bm{\Lambda}_{n}\right]\leq C_{K}(\textsf{E}Z^{2})^{1/2}\frac{\|\bm{\Lambda}_{n}\|^{K}}{n^{K\gamma}}.$
Choose $K>2$ such that $K\gamma/2>1$. By (2.9),
@@
The proof is completed by (4.1) and (2.9). $\Box$
Proof of Theorem 2.4. For (i), notice that (4.1) still holds for $\gamma=0$. When $\|\bm{\Lambda}_{n}\|\geq\frac{2c_{r}\beta_{2}}{r\lambda_{1}\lambda_{2}}+1$,
$\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}\leq-\frac{r\lambda_{1}\lambda_{2}}{2}\|\bm{\Lambda}_{n}\|^{r-1}+c_{r}\beta_{r}.$
When $\|\bm{\Lambda}_{n}\|\leq\frac{2c_{r}\beta_{2}}{r\lambda_{1}\lambda_{2}}+1$, $\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}\leq-\frac{r\lambda_{1}\lambda_{2}}{2}\|\bm{\Lambda}_{n}\|^{r-1}+C_{r}$. Thus, exist $a_{r},b_{r}>0$ such that $\textsf{E}\left[\|\bm{\Lambda}_{n+1}\|^{r}|\mathscr{F}_{n}\right]-\|\bm{\Lambda}_{n}\|^{r}\leq-a_{r}\|\bm{\Lambda}_{n}\|^{r-1}+b_{r}.$ It follows that
$\textsf{E}[\|\bm{\Lambda}_{n+1}\|^{r}]-\textsf{E}[\|\bm{\Lambda}_{n}\|^{r}]\leq$ $-a_{r}\textsf{E}[\|\bm{\Lambda}_{n}\|^{r-1}]+b_{r}$ (4.43)
$\leq$ $b_{r}.$ (4.44)
From (4.44), it follows that $\textsf{E}[\|\bm{\Lambda}_{n}\|^{r_{0}}]\leq nb_{r_{0}}$.
Now, let $m:=m_{n}\in\{0,1,\ldots,n\}$ be the last one for which $a_{r}\textsf{E}[\|\bm{\Lambda}_{m}\|^{r-1}]+b_{r}\geq 0.$ Let $r-1+q=r_{0}$, $1/p+1/q=1$. Then
$\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\leq$ $\textsf{E}\left[\|\bm{\Lambda}_{m+1}\|^{r}\right]\leq\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r}\right]+b_{r}=b_{r}+\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{\frac{r-1}{p}}\|\bm{\Lambda}_{m}\|^{\frac{r-1}{q}+1}\right]$
$\leq$ $b_{r}+\left(\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-1}\right]\right)^{1/p}\left(\textsf{E}\left[\|\bm{\Lambda}_{m}\|^{r-1+q}\right]\right)^{1/q}$
$\leq$ $b_{r}+(b_{r}/a_{r})^{1/p}\left(b_{r_{0}}m\right)^{1/q}\leq C_{r,r_{0}}n^{1/q}.$ (4.45)
Letting $r=2$ and $q=r_{0}-1$ yields $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{2}\right]=O(n^{\frac{1}{r_{0}-1}})$. When $r_{0}\geq 3$. Let $r=(1+r_{0})/2$. Then $r\geq 2$, $q=r$. Thus, $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]=O(n^{1/r})$, and so $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{2}\right]=O(n^{2/r^{2}})=O(n^{\frac{8}{(r_{0}+1)^{2}}}$. The proof of (i) is completed.
@@ 4.2 The properties of the test
$\widehat{\tau}_{n}=:$ $\overline{Y}_{n,1}-\overline{Y}_{n,1}=\tau+\frac{\sum_{i=1}^{n}e_{i,1}+\frac{D{{}_{n,1}}}{\rho}}{n\big(1+\frac{D_{n}}{n\rho}\big)}-\frac{\sum_{i=1}^{n}e_{i,2}-\frac{D{{}_{n,2}}}{1-\rho}}{n\big(1-\frac{D_{n}}{n(1-\rho)}\big)}$
$=$ $\tau+\frac{\frac{D{{}_{n,1}}}{\rho}+\frac{D{{}_{n,2}}}{1-\rho}+\sum_{i=1}^{n}(e_{i,1}-e_{i,2})}{n}+o_{P}(n^{1/2})$
$=$ $\tau+\frac{\sum_{i=1}^{n}(T_{i}-\rho)\big(\frac{e{{}_{i,1}}}{\rho}+\frac{e{{}_{i,2}}}{1-\rho}\big)+\sum_{i=1}^{n}(e_{i,1}-e_{i,2})}{n}+o_{P}(n^{1/2}).$
@@
$=$ $\frac{\textsf{E}\tilde{e}_{1}^{2}+\rho(1-\rho)\textsf{E}(\textsf{Pj}[e_{1}-e_{2}|\bm{\phi}])^{2}}{\rho}+\frac{\textsf{E}\tilde{e}_{2}^{2}+\rho(1-\rho)\textsf{E}(\textsf{Pj}[e_{1}-e_{2}|\bm{\phi}])^{2}}{1-\rho}.$

### v3

@@ 4.1 The properties of the CAR procedures
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+O\Big(n^{\frac{\gamma}{2}}+n^{\frac{(r_{0}-2)\gamma+2}{4(r_{0}-1)}}+n^{1-\gamma}+n^{1+\frac{(r_{0}-2)\gamma+2}{2(r_{0}-1)}-2\gamma}\Big)$
$=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+O\Big(n^{\frac{\gamma}{2}}+n^{1-\gamma}\Big)\text{ in }L_{1}$
by the fact that $\frac{(r_{0}-2)\gamma+2}{2(r_{0}-1)}\leq\gamma$ when $\gamma\geq 2/r_{0}$. ((iii)) is proven.
(iv) For (2.11), notice that $\{(\Delta M_{i,\tilde{Z}},W_{i}),\mathscr{F}_{i},i=1,2,\ldots\}$ is sequence of martingale differences. We have $|\Delta M_{i,\tilde{Z}}|\leq|\widetilde{Z}_{i}|+\textsf{E}|\widetilde{Z}_{i}|$,
@@
i.e., the Linderberg condition is satisfied. Further, $\textsf{E}[W_{i}^{2}|\mathscr{F}_{i-1}]=\textsf{E}W^{2}$,
@@
$\left(\frac{M_{n,\widetilde{Z}}}{\sqrt{n}},\frac{\sum_{i=1}^{n}W_{i}}{\sqrt{n}}\right)\overset{d}{\to}N(0,\vec{\sigma}^{2})\times N(0,\textsf{Var}(W)),$
by the central limit theorem for martingales(Hall and Heyde, , 1980, Corollary 3.1). Hence, (2.11) holds by ((iii)) and the fact that $1/2<\gamma<1$.
When $(r_{0}-2)\gamma\geq 2$, applying (2.5) with $r=4$ yields that
$\|\bm{\Lambda}_{n}\|=O\big(n^{\gamma/2}\big)\text{ in }L_{4}.$
Thus, ((iii)) holds in $L_{2}$ by (4.42). Then, (2.12) holds by (4.43). $\Box$
Proof of Remark 2.2. Notice $\big|\ell(x)-\rho-\ell^{\prime}(0)x-\frac{1}{2}\ell^{\prime\prime}(0)x^{2}\big|\leq c|x|^{3}.$ Similar to (4.1), we have
@@
$\leq$ $C\textsf{E}\left[\frac{\|\bm{\Lambda}_{n}\|^{3}\|\bm{\phi}\|^{3}|\widetilde{Z}|}{n^{2\gamma}}\big|\bm{\Lambda}_{n}\right]\leq C\frac{\|\bm{\Lambda}_{n}\|^{3}}{n^{2\gamma}}\big(\textsf{E}\|\bm{\phi}\|^{6}\big)^{1/2}\big(\textsf{E}\widetilde{Z}^{2}\big)^{1/2}\leq C_{0}(\textsf{E}Z^{2})^{1/2}\frac{\|\bm{\Lambda}_{n}\|^{3}}{n^{3\gamma}}.$
@@
$=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+O\Big(n^{\gamma/2}+n^{\frac{\gamma}{2}-\frac{(r_{0}-1)\gamma-2}{2(r_{0}-2)}}+n^{1-3\gamma/2}+n^{1-\frac{3\gamma}{2}-\frac{(r_{0}-1)\gamma-2}{6(r_{0}-2)}}\Big)\text{ in }L_{1}$
$=$ $\sum_{i=1}^{n}\Delta M_{i,\widetilde{Z}}+o(n^{1/2})\text{ in }L_{1},$
by (2.5) (with $r=3$) and the facts that $\gamma/2<1$, $1-\frac{3\gamma}{2}<1/2$ and $(r_{0}-1)\gamma-2\geq 0$ when $1/3<\gamma<1$ and $r_{0}\geq 7$, where $C_{0}$ does not depend on $Z$. $\Box$.
Proof of Theorem 2.2. Notice $\Phi(u_{\rho/2})<\rho<\Phi(u_{(\rho+1)/2})$. Let $K\geq 2$. By the continuity, there exists $\delta>0$ such that
@@
When $|x|\geq\delta$, we have $|\ell(x)-\rho|\leq 1\leq\delta^{-K}|x|^{K}$, $|x|\leq\delta^{-K+1}|x|^{K}$. So, there exists a constant $C_{K}$ such that
@@
$\;\;\leq c_{K}\textsf{E}\left[\frac{\|\bm{\Lambda}_{n}\|^{K}\|\bm{\phi}\|^{K}|Z|}{n^{2\gamma}}\big|\bm{\Lambda}_{n}\right]\leq c_{K}(\textsf{E}Z^{2})^{1/2}(\textsf{E}\|\bm{\phi}\|^{2K})^{1/2}\frac{\|\bm{\Lambda}_{n}\|^{K}}{n^{K\gamma}}.$ (4.44)
Now, let $K=r_{0}/2$. By (2.7), $\|\bm{\Lambda}_{n}\|=o(n^{1/2})$ in $L_{2}$, and
$\frac{\|\bm{\Lambda}_{n}\|^{K}}{n^{K\gamma}}=O\left(\frac{n^{r_{0}\gamma/4+(1-\gamma)/2}}{n^{r_{0}\gamma/2}}\right)=O\left(n^{\frac{1}{2}-\frac{(r_{0}+2)\gamma}{4}}\right)=o(n^{-1/2})\text{ in }L_{2}.$
Thus, by (4.1),
$\sum_{i=1}^{n}(T_{i}-\rho)Z_{i}=M_{n,\widetilde{Z}}+o(n^{1/2})\text{ in }L_{2},$
which, together with (4.43), implies (2.11) and (2.12). $\Box$
Proof of Theorem 2.3. (i). For fixed $r\geq 2$ and $0<\gamma<1$, choose $r_{0}>2$ large enough such that $(r_{0}+2-r)\gamma-2\geq 0$. Then by (2.5),
$\left(\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]\right)^{1/r}\leq C_{r,r_{0}}n^{\frac{\gamma}{2}}.$
(2.14) is proven.
(ii) Notice (4.1). Choose $K>2$ such that $K\gamma/2>1$. By (2.14),
@@
The proof of (2.15) is completed by (4.1) and (2.14).
(2.11) and (2.12) follow from (2.15) and (4.43). $\Box$
Proof of Theorem 2.4. For (i), notice that (4.1) and (4.38) still hold for $\gamma=0$. Thus, (2.5) and (2.6) still hold for $\gamma=0$. Letting $r=2$ and $g=r_{0}-1$ in (2.5) yields $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{2}\right]=O(n^{\frac{1}{r_{0}-1}})$. When $r_{0}\geq 3$. Let $r=(1+r_{0})/2$ in (2.5). Then $r\geq 2$, $g=r$. Thus, $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{r}\right]=O(n^{1/r})$, and so $\textsf{E}\left[\|\bm{\Lambda}_{n}\|^{2}\right]=O(n^{2/r^{2}})=O(n^{\frac{8}{(r_{0}+1)^{2}}})$. The proof of (i) is completed.
@@ 4.2 The properties of the test
$\widehat{\tau}_{n}=:$ $\overline{Y}_{n,1}-\overline{Y}_{n,2}=\tau+\frac{\sum_{i=1}^{n}e_{i,1}+\frac{D{{}_{n,1}}}{\rho}}{n\big(1+\frac{D_{n}}{n\rho}\big)}-\frac{\sum_{i=1}^{n}e_{i,2}-\frac{D{{}_{n,2}}}{1-\rho}}{n\big(1-\frac{D_{n}}{n(1-\rho)}\big)}$
$=$ $\tau+\frac{\frac{D{{}_{n,1}}}{\rho}+\frac{D{{}_{n,2}}}{1-\rho}+\sum_{i=1}^{n}(e_{i,1}-e_{i,2})}{n}+o_{P}(n^{-1/2})$
$=$ $\tau+\frac{\sum_{i=1}^{n}(T_{i}-\rho)\big(\frac{e{{}_{i,1}}}{\rho}+\frac{e{{}_{i,2}}}{1-\rho}\big)+\sum_{i=1}^{n}(e_{i,1}-e_{i,2})}{n}+o_{P}(n^{-1/2}).$
@@
$=$ $\frac{\textsf{E}\tilde{e}_{1}^{2}}{\rho}+\frac{\textsf{E}\tilde{e}_{2}^{2}}{1-\rho}+\textsf{E}(\textsf{Pj}[e_{1}-e_{2}|\bm{\phi}])^{2}.$